A city farm is rectangular in shape. It is 65 meters long and 60 meters wide.
Calculate the length of the perimeter.

Answer by macston(5194)

.
L=length=65m; W=width=60m; P=perimeter
P=2(L+W)
P=2(65m+60m)
P=2(125m)
P=250m
ANSWER: The perimeter is 250 meters.